OpenStudy (anonymous):

find the limit of: Limit(x->0) : (cosx-1)/sinx All i know is you start by performing the quotient rule, but now im stuck! thanks in advance!

zepdrix (zepdrix):

Have you learned `L'Hospital's Rule` at this point?

zepdrix (zepdrix):

\[\Large\bf\sf \lim_{x\to0}\frac{\cos x-1}{\sin x}\] We're getting the indeterminate form:\[\Large\bf\sf \frac{0}{0}\]So we have a good candidate for L'H Rule. But have you learned that method yet?

OpenStudy (anonymous):

i have not...not yet anyways

zepdrix (zepdrix):

Hmm ok here's another method we can try. Let's multiply the top and bottom by the `conjugate` of the top.

zepdrix (zepdrix):

\[\Large\bf\sf \lim_{x\to0}\frac{\cos x-1}{\sin x}\left(\frac{\cos x+1}{\cos x+1}\right)\quad=\quad \lim_{x\to0}\frac{\cos^2-1}{\sin x(\cos x+1)}\]

zepdrix (zepdrix):

If you factor a -1 out of the top, you can apply your square identity.\[\Large\bf\sf \lim_{x\to0}\frac{-(1-\cos^2x)}{\sin x(\cos x+1)}\quad=\quad \lim_{x\to0}\frac{-(\sin^2x)}{\sin x(\cos x+1)}\]

zepdrix (zepdrix):

From here we can cancel out some sines, yes? Any step in there really confusing?

zepdrix (zepdrix):

This is something to try and get used to with limits. You plug the value directly in, if there is a problem, you do some algebra and try to cancel stuff out. After canceling something out, see if you still have a problem plugging the value directly in.

OpenStudy (anonymous):

thank you sooo much!!! this helped tons!!!!! i understand! wahoo

zepdrix (zepdrix):

Oh good c: